As a healthcare executive, you have two options; One procedure has a cost of $300,000, and increases QALY (Quality-Adjusted Life Year) by 8, another is $180,000 and a QALY of 5. Which one would you choose and why.
Answer #1

ANSWER:

We need to find the cost per qaly gained.

1st procedure cost per qaly gained = cost / qaly = 300,000 / 8 = 37,500

2nd procedure cost per qaly gained = cost / qaly = 180,000 / 5 = 36,000

so the 2nd procedure has a lower cost per qaly gained and so therefore we will choose the 2nd procedure.